# Video-Driven Graph Network-Based Simulators

### v2

@@ 4 Results
Finally, we analyze the interpolation capabilities of our model. While the system classes lack interoperability, we exploit the implementation of the simulator. The sand class has a parameter for the angle of friction, set to 45 degrees by default. When set to 0, the system behaves like the water class. To test the model’s ability to interpolate between systems, we create an additional evaluation dataset. This dataset includes 4 trajectories for 15 system classes, with the friction angle increasing from 0 to 45 degrees. We then evaluate the system’s one-step accuracy on this dataset. The bottom-right plot in Figure 2 shows the one-step errors for varying angles of friction.
@@ Appendix B Model Architecture
The Video Encoder processes a series of RGB images. Each frame is encoded into a low-dimensional representation with a series of linear layers, each followed by a tanh activation function. This output serves as the input for an LSTM with a 32-dimensional hidden state. Finally, the LSTM output is mapped to a 4-dimensional physical encoding.
Functions constituting a Graph Network are implemented as multi-layer perceptions with a single hidden layer and a tanh activation. Table 2 summarizes the model’s architecture and exact dimensionality.

### v3

@@ 4 Results
Finally, we analyze the interpolation capabilities of our model. While the system classes lack interoperability, we exploit the implementation of the simulator. The sand
class has a parameter for the angle of friction, set to 45 degrees by default. When set to 0, the system behaves like the water
class. To test the model’s ability to interpolate between systems, we create an additional evaluation dataset. This dataset includes 4 trajectories for 15 system classes, with the friction angle increasing from 0 to 45 degrees. We then evaluate the system’s one-step accuracy on this dataset. The bottom-right plot in Figure 2 shows the one-step errors for varying angles of friction.
@@ Appendix B Model Architecture
The Video Encoder processes a series of RGB images. Each frame is encoded into a low-dimensional representation with a series of linear layers, each followed by a tanh
activation function. This output serves as the input for an LSTM with a 32-dimensional hidden state. Finally, the LSTM output is mapped to a 4-dimensional physical encoding.
Functions constituting a Graph Network are implemented as multi-layer perceptions with a single hidden layer and a tanh
activation. Table 2 summarizes the model’s architecture and exact dimensionality.